OpenStudy (anonymous):

NEED HELP please: Cos (sin^-1(1/3) - tan^-1(1/2)) Find the exact value of the expression.

OpenStudy (anonymous):

0.992345241211553801626983189764143358021255927794159535446727...

OpenStudy (anonymous):

i need the answer in simplest fraction form, thanks tho

OpenStudy (anonymous):

(4 sqrt(2/5))/3+1/(3 sqrt(5))

OpenStudy (anonymous):

The textbook answer is (4sqrt10 + sqrt5)/15 But i need to see how you do it.

OpenStudy (anonymous):

alternate form=1/15 (4 sqrt(10)+sqrt(5))

OpenStudy (anonymous):

Can you explain to me how you got that?

OpenStudy (anonymous):

cos(sin^(-1)(1/3)-tan^(-1)(1/2)) (result in radians)

OpenStudy (anonymous):

cos(cot^(-1)(2)-csc^(-1)(3))

OpenStudy (anonymous):

cos(sin^(-1)(1/3)-tan^(-1)(1/2)) = cosh(i (sin^(-1)(1/3)-tan^(-1)(1/2)))

OpenStudy (anonymous):

cos(sin^(-1)(1/3)-tan^(-1)(1/2)) = cosh(-i (sin^(-1)(1/3)-tan^(-1)(1/2)))

OpenStudy (anonymous):

cos(sin^(-1)(1/3)-tan^(-1)(1/2)) = 1/2 (e^(-i (sin^(-1)(1/3)-tan^(-1)(1/2)))+e^(i (sin^(-1)(1/3)-tan^(-1)(1/2))))

OpenStudy (anonymous):

cos(tan^(-1)(1/2)-tan^(-1)(1/(2 sqrt(2))))

OpenStudy (anonymous):

(4 sqrt(2/5))/3+1/(3 sqrt(5))

OpenStudy (anonymous):

alternate form=1/15 (4 sqrt(10)+sqrt(5))

OpenStudy (anonymous):

alright thank you very much

OpenStudy (anonymous):

Do I look dumb?

OpenStudy (anonymous):

What do you mean?

OpenStudy (anonymous):

you dont trust me?

OpenStudy (anonymous):

haha i had no doubt you had the answer right but i trying to review, so i wanted to see work to help me.

OpenStudy (anonymous):

once again, thanks for your help wiz :)